# Under-Canopy Navigation using Aerial Lidar Maps

### v1

@@ V Global Path Planning
Our path planning module plans a 2D path (over the ground cells) for the ground robot given a pair of start and goal locations. A path $\mathcal{P}$ is a sequence of adjacent ground cells $\mathcal{P}=(s_{0},s_{1},\ldots,s_{\ell})$ starting from the current cell $s_{0}$ and ending at the goal cell $s_{\ell}=s_{\text{goal}}$. The cost of transitioning from cell $s$ to one of its adjacent cells $s^{\prime}$ (hereafter, the edge cost) is denoted by $c(s,s^{\prime})\in\mathbb{R}^{+}$. We use D* Lite [3] to find a minimum-cost path where the path cost is given by the sum of its edge costs,
$J({\mathcal{P}})\triangleq\sum_{i=0}^{\ell-1}c(s_{i},s_{i+1}).$ (4)
D* Lite allows us to efficiently replan an optimal path from the current cell to the goal based on the latest edge costs upon encountering a blockage. In the following, we introduce two choices used in our simulations and experiments for the edge costs $c(s,s^{\prime})$ that incorporate both the footprint obstruction score $b_{\text{max}}(s)$ (reflecting the risk of choosing a blocked path) and the travelled distance.
@@ V-1 Expected Cost
This edge cost function is inspired by the expected traversal cost proposed in [27, Eq. 8]. Formally, the edge cost between two adjacent cells $s_{i}$ and $s_{i+1}$ in a given path is defined as follows,
@@ V-2 Log-Reachability cost
Our second path cost is based on the reachability metric proposed in [28]. Informally speaking, reachability of a path reflects the likelihood of reaching the goal cell without hitting any blocked cells. We approximate the product integral derived in [28, Eq. 14] for computing the reachability of the goal-directed path $\mathcal{P}$ by the following expression,
@@
where, as before, $d(s,s^{\prime})$ denotes the octile distance between $s$ and $s^{\prime}$, and $b_{\text{min}}$ is the minimum obstruction score across the entire map.
@@ VI Simulations
We validated our proposed approach using the Gazebo simulator [^2] https://gazebosim.org/. We created three synthetic forests using vegetation models of the NEGS-UGV Dataset [29], and acquired simulated airborne lidar scans of these forest worlds. In the following, we first evaluate the accuracy of our proposed expected occupancy mapping approach, and then analyse the performance of our full navigation system.
VI-A Ablation Study: Expected Occupancy Mapping
@@
The objective is to evaluate the accuracy of our expected occupancy map (with localisation uncertainty) and compare it with the standard occupancy map for different levels of error in the sensor trajectory. We created three random forest worlds named Forest I, II and III, and simulated three different aerial trajectories (one per scenario), as shown in Fig. 4, of a VLP-16 lidar sensor flying above the canopy and scanning the forests. We constructed ground-truth occupancy maps of each scenario by simulating an ideal sensor (no noise in range measurements or trajectory) and processing the ground-truth point clouds and trajectories using OHM. Then, we perturbed each sensor pose along the trajectories with independent, additive Gaussian noise $\sim\mathcal{N}\left(0,\sigma^{2}\right)$ on each component to simulate a (perturbed) localisation estimate of the sensor trajectory. For each perturbation level, specified by standard deviation values, we generated 10 perturbed trajectories and point clouds for subsequent comparisons. We also added Gaussian noise with a fixed standard deviation ($\sigma_{r}=0.01\,$m) to range measurements to simulate sensor uncertainty. For our expected occupancy mapping (Sec. IV-B) we chose $N=20$.
@@
Figure 4: (a): Example of synthetic forest created in Gazebo where we simulated lidar sensors moving above the canopy to evaluate our method for incorporation of uncertainty from sensor poses to occupancy values. (b) to (d): Ground-truth trajectories of the airborne sensor and the corresponding registered scans for the datasets acquired in the synthetic environments Forest I, II and III, respectively. After adding perturbations to sensor poses, a perturbed point cloud of each dataset was produced to be processed by our approach.
@@
With full knowledge of ground-truth occupancy probabilities from the ideal sensor, and the perturbed datasets, we can further inspect the maps’ accuracy using the Kullback-Leibler Divergence (KLD) of voxel occupancy probabilities between the expected and standard occupancy maps with respect to the ground-truth map.
@@
In Forest I, we analysed the isolated effect of perturbations in sensor’s translation and orientation, respectively. Since the percentage of occupied voxels (about $10\%$) in the ground-truth map is considerably lower than the free counterparts (about $80\%$), we also computed the average KLD of expected and standard occupancy maps per voxel class (free, occupied and unknown) in this first scenario. Ground-truth probability ($p_{\text{gt}}$) thresholds are used to classify voxels as free ($p_{\text{gt}}<0.45$), unknown ($0.45\leq p_{\text{gt}}\leq 0.55$) or occupied ($p_{\text{gt}}>0.55$).
@@
Figure 5: Comparison of the expected occupancy against the standard occupancy map for different perturbation values added to sensor poses. (a) and (b) show the average KLD values in Forest I dataset for perturbations in position only and orientation only, respectively. KLD values are depicted per voxel class on the ground-truth dataset. (c) and (d) illustrate the KLD for all voxels classes in datasets Forest II and Forest III, respectively, for combined perturbation levels (position and orientation simultaneously) where $\sigma=(0.015\text{m},0.015\text{m},0.015\text{m},0.11^{\circ},0.11^{\circ},0.11^{\circ})$. In the figures, noise refers to perturbations added to sensor poses.
@@
Results illustrated in Fig. 5 (a) and (b) show that for both perturbations (position and orientation), the maps generated using expected occupancy presented lower average KLD for all perturbation levels, and that this behaviour was consistent among all voxels classes, confirming that our method improves the quality of maps on occupied, free and unknown regions. We also assessed the effect of combined perturbations (position and orientation jointly) on map quality in Forest II and III (Fig. 5 (c) and (d), respectively). As observed, the average KLD (all voxels) on the expected occupancy also outperforms the standard occupancy with combined perturbation values.
@@
Finally, we computed the Area Under Curve (AUC) of the Receiver Operating Characteristic (ROC) curves for the standard occupancy and expected occupancy approaches, providing an overview of each method’s performance as predictors of free or occupied states in datasets Forest II and III. Results are summarised in Table I, where AUC values closer to one indicate better prediction rates. For increasing levels of perturbation, both approaches experience reduction of AUC, which is anticipated since higher errors on sensor poses negatively affect the quality of occupancy values in both methods. However, in both datasets, the expected occupancy method displayed higher AUC than the standard occupancy, showing the benefits of incorporating localisation uncertainty in the occupancy calculation.
@@
TABLE I: Comparison of the AUC values for the standard occupancy (occ) and expected occupancy (exp) approaches for increasing perturbation levels in sensor poses, where $\sigma=(0.015\text{m},0.015\text{m},0.015\text{m},0.11^{\circ},0.11^{\circ},0.11^{\circ})$. Higher AUC values indicate better map predictions.
Forest II Forest III
Perturbation level occ exp occ exp
$\sigma$ 0.982 0.988 0.984 0.990
@@ VI-B Full system evaluation
To assess whether our navigation approach (mapping and planning), produces more efficient paths in comparison with a naive baseline planner without prior maps, we also created a simulated forest with ground obstacles as illustrated in Fig. 6. Similarly to Sec. VI-A, we generated a lidar trajectory above the canopy with poses perturbed using $\sigma=(0.05\text{m},0.05\text{m},0.05\text{m},0.11^{\circ},0.11^{\circ},0.11^{\circ})$. We then processed the lidar scans using our mapping method (Sec. IV) to create the obstruction map used by the D* Lite planner.
Figure 6: (a) Gazebo forest model used in simulations of the planning methods. (b) Shows the density of ground obstacles (red) composed by high vegetation and tree trunks in the synthetic forest.
Figure 7: Executed paths comparison between our planner using three cost functions: log-reachability cost, expectation cost with ($C_{\text{obst}}:5$) and ($C_{\text{obst}}:20$), against the naive planner using only local information. (a) shows the histogram of path length ratio between the executed path of our method and the naive approach. (b) depicts the cumulative count of paths for the same ratios. Curves closer to the top left show better performance.
@@
Figure 8: The satellite image (a) and the aggregated lidar scans (b) of the forest area used in our experiments. The resulting obstruction map generated using our proposed approach is shown in (c) along with a magnified section for better visualisation in (d).
@@
We defined 40 random start and end goal pairs in the synthetic forest to be executed by our navigation method and the naive planner. The naive planner does not use the prior map, and is equivalent to assuming a direct path from the start to the goal as the global prior path. All 40 runs were repeated with three configurations of cost function: log-reachability cost, and two versions of expected cost (5) with values $C_{\text{obst}}=5$ and $C_{\text{obst}}=20$, respectively. During online execution of the planners, the ground-truth is utilised to update the obstruction map in our navigation approach and to construct a map on the fly in the naive baseline. We defined a 5 meter radius around the robot pose to update the maps with local information from the ground-truth environment, hence map cells in this region are changed to near 1 or 0, in case obstacles or free areas are detected, respectively, in the same cell coordinates of the ground-truth map. In both methods we utilise the D* Lite planner to replan in case obstacles block the robot’s planned path. While our method replans on a combined map (prior map with local updates), the naive approach replans only on the map built online with local information.
The metric used for evaluation is the ratio of the executed path length between our navigation method and the naive baseline. Fig. 7 (a) and (b) show the results of our complete navigation framework using the chosen configurations of cost function. As observed in the histogram, using the log-reachability cost resulted in $87.5\%$ of executed paths having lengths shorter than the naive planner, confirming the clear advantage of using our mapping method as prior information for global planning and also online execution and replanning. This result also confirms that even in the presence of perturbations in the sensor trajectory and partial canopy cover, the proposed expected occupancy and obstruction mapping approaches provide enough information for generating efficient paths.
The expectation cost, however, resulted in $47.5\%$ of paths shorter than the baseline. This performance reduction can be attributed to the fact that when using the expected cost some planned paths crossed a small number of cells with high obstruction scores instead of planning longer detours. During execution with the ground robot, some of these paths tended to guide the robot to blocked areas in the ground-truth map (in some instances multiple times), requiring expensive local replanning steps. Further, although the naive approach has no prior, in cases with few obstructions the naive assumption (straight-line) may actually be globally optimal or near-optimal, requiring little or no replanning. Because the log-reachability cost is more risk averse by design, it may prefer to detour even small potentially obstructed areas, yielding better results (see Fig. 7 (b)).
@@ VII Real-World Experiments
Figure 9: Results of real world experiments Test 1 (a) and Test 2 (b) showing the executed paths of our approach and the naive method. Black circle and star represent the start and goal locations, respectively. In both experiments our approach guided the robot to shorter and less cluttered paths, whereas the limited look-ahead of the naive method resulted in a longer trajectory in (a) due to crossing a more cluttered area, and ultimate failure of the planner in (b) by facing a dead-end of high undergrowth.
@@
Since it is not trivial to define ground-truth for obstructed areas in forest environments given their complexity, e.g., cells that are visually “obstructed” by vegetation are often traversable, we analysed qualitatively our mapping output from real data. As observed in Fig. 8 (a)-(d), even with partial canopy occlusion, our mapping using above-canopy lidar scans is capable of detecting regions at ground level with high obstruction scores (yellow to red) indicating the presence of obstacles, and free regions (green) with higher chances of traversing. Uncertain areas (blue) are also captured, providing meaningful information for planning and decision-making.
@@
To test our full system, we defined two start and end goal pairs in the forest named Test 1 and 2. For each test location we performed two runs, one with the naive approach where the robot used only the local perception and its embedded navigation system to reach the target without any prior information, and another run using our proposed navigation method (prior aerial mapping and global planning) with the log-reachability cost as it demonstrated superior performance in simulation. Unlike the navigation pipeline used in simulations, here our method was used only offline to create the obstruction map and plan a global trajectory from start to goal using D* Lite. Then, for online execution, the only prior information provided to the robot was the planned trajectory as a set of waypoints followed by the local navigation system [23] of the DTR. The local planner (Hybrid A* [30]) and perception allow the robot to deviate around locally-detected obstacles while tracking the waypoints. For fair comparison, the same local navigation system was used for both runs, naive and our method; in this manner, we can assess the benefits of providing information prior to operation through a global trajectory.
@@
During Test 1, depicted in Fig. 9 (a), while our approach preferred a less cluttered path on the left side of the large obstacle, the naive approach took the right side, following a much denser path that also crossed a narrow passage among high bushes. Our method travelled a path length of $69.6$ m in comparison to $79.9$ m of the naive trajectory, which demonstrates that our proposed navigation not only increases path efficiency, but it is also more risk averse preferring safer areas with less likelihood of having obstacles.
In Test 2, Fig. 9 (b), the limited look-ahead of the naive approach directed the robot to a dead-end composed of cluttered high vegetation. The local navigation was unable to detour the blocked area, and failed to reach the goal. Unlike the naive counterpart, our method leveraged the prior information, avoiding the dead-end and taking a clear path on the right of the vegetated area to reach the goal.
@@
Although we acknowledge that the practical experiments do not represent statistically representative results, the outcomes validate the beneficial behaviours of our method observed in simulation. The field trials showed that planning using our obstruction maps in tandem with the proposed log-reachability cost leads to shorter and safer paths by avoiding cluttered areas, which ultimately increases the mission success rate.

### v2

@@ V Global Path Planning
Our path planning module plans a 2D path (over the ground cells) for the ground robot given a pair of start and goal locations. A path $\mathcal{P}$ is a sequence of adjacent ground cells $\mathcal{P}=(s_{0},s_{1},\ldots,s_{\ell})$ starting from the current cell $s_{0}$ and ending at the goal cell $s_{\ell}=s_{\text{goal}}$. The cost of transitioning from cell $s$ to one of its adjacent cells $s^{\prime}$ (hereafter, the edge cost) is denoted by $c(s,s^{\prime})\in\mathbb{R}^{+}$. We use D* Lite [3] to find a minimum-cost path where the path cost is given by the sum of its edge costs $J({\mathcal{P}})\triangleq\sum_{i=0}^{\ell-1}c(s_{i},s_{i+1})$. D* Lite allows us to efficiently replan an optimal path from the current cell to the goal based on the latest edge costs upon encountering a blockage. In the following, we introduce two choices used in our simulations and experiments for the edge costs $c(s,s^{\prime})$ that incorporate both the footprint obstruction score $b_{\text{max}}(s)$ (reflecting the risk of choosing a blocked path) and the travelled distance.
@@ V-1 Expected Cost
This edge cost function is inspired by the expected traversal cost proposed in [25, Eq. 8]. Formally, the edge cost between two adjacent cells $s_{i}$ and $s_{i+1}$ in a given path is defined as follows,
@@ V-2 Log-Reachability cost
This path cost is based on the reachability metric proposed in [26]. Informally speaking, reachability of a path reflects the likelihood of reaching the goal cell without hitting any blocked cells. We approximate the product integral derived in [26, Eq. 14] for computing the reachability of the goal path $\mathcal{P}$ by the following expression,
@@
where, as before, $d(s,s^{\prime})$ is the octile distance between $s$ and $s^{\prime}$, and $b_{\text{min}}$ is the minimum obstruction score in the map.
@@ VI Simulations
We validated our proposed approach using the Gazebo simulator [^2] https://gazebosim.org/. We created synthetic forests using vegetation models of the NEGS-UGV Dataset [27], and acquired simulated aerial lidar scans of these forest worlds. In the following, we first evaluate the accuracy of our proposed UA-occupancy mapping, and then analyse the performance of our full navigation system.
VI-A Ablation Study: UA-Occupancy Mapping
@@
First, we evaluate the accuracy of our UA-occupancy map (with pose uncertainty) and compare it with the standard occupancy map for different levels of error in the sensor trajectory. We created three random forest worlds: Forest I, II and III, and simulated three different aerial trajectories (one per scenario), as shown in Fig. 4, of a VLP-16 lidar sensor flying above the canopy and scanning the forests. We also simulated trajectory A in Forest II and III. We constructed ground-truth occupancy maps for each scenario by simulating an ideal sensor (no noise in range measurements or trajectory) and processing the ground-truth point clouds and trajectories using OHM. Then, we perturbed each sensor pose along the trajectories with independent, additive Gaussian noise $\sim\mathcal{N}\left(0,\sigma^{2}\right)$ on each component $[x,y,z,\text{roll},\text{pitch},\text{yaw}]$ to simulate a GPS-like estimate of the sensor trajectory. Therefore, perturbations are not accumulated along the trajectory. For each perturbation level, specified by standard deviation values, we generated 10 perturbed trajectories and point clouds for subsequent comparisons. We added Gaussian noise with a fixed standard deviation ($\sigma_{r}=0.01\,$m) to range measurements to simulate sensor uncertainty. For our UA-occupancy mapping (Sec. IV-B) we chose $N=20$.
@@
Figure 4: (a): Example of synthetic forest created in Gazebo where we simulated lidar scans captured above the canopy. (b) to (d): Ground-truth trajectories (A, B and C) of the airborne sensor and the corresponding registered scans for the datasets acquired in the synthetic environments Forest I, II and III, respectively. After adding perturbations to sensor poses, a perturbed point cloud of each dataset was produced to be processed by our approach.
@@
With full knowledge of ground-truth occupancy probabilities from the ideal sensor, and the perturbed datasets, we can further evaluate map accuracy using the Kullback-Leibler Divergence (KLD) of voxel occupancy values between our UA and standard occupancy maps with respect to the ground-truth.
@@
In Forest I, we analysed the isolated effect of translation and orientation perturbations. Since the percentage of occupied voxels (about $10\%$) in the ground-truth map is considerably lower than the free counterparts (about $80\%$), we also computed the average KLD of our UA and standard occupancy maps per voxel class (free, occupied and uncertain) in this first scenario. Ground-truth probability ($p_{\text{gt}}$) thresholds are used to classify voxels as free ($p_{\text{gt}}<0.45$), uncertain ($0.45\leq p_{\text{gt}}\leq 0.55$) or occupied ($p_{\text{gt}}>0.55$).
@@
Figure 5: Comparison of our UA-occupancy against the standard occupancy map for different perturbation values added to sensor poses. (a) and (b) show the average KLD values in Forest I (trajectory A) for perturbations in position only and orientation only, respectively. KLD values are depicted per voxel class on the ground-truth dataset. (c) and (d) illustrate the KLD values for all voxel classes in Forest II (2 trajectories) and Forest III (2 trajectories), respectively, for combined perturbation levels (position and orientation) where $\sigma=(0.015\text{m},0.015\text{m},0.015\text{m},0.11^{\circ},0.11^{\circ},0.11^{\circ})$. In the figures, noise refers to perturbations added to sensor poses and error bars show $\pm$ std. deviation.
@@
Results illustrated in Fig. 5 (a) and (b) show that for both perturbations (position and orientation), the maps generated using our UA-occupancy presented lower average KLD for all perturbation levels, and that this behaviour was consistent among all voxel classes, confirming that our method improves the quality of maps on occupied, free and uncertain regions. We also assessed the effect of combined perturbations (position and orientation jointly) on map quality in Forest II and III (Fig. 5 (c) and (d), respectively). As observed, the average KLD (all voxels) on the UA-occupancy also outperforms the standard occupancy with combined perturbation values.
@@
Finally, we computed the Area Under Curve (AUC) of the Receiver Operating Characteristic (ROC) curves for the standard occupancy and UA-occupancy approaches, providing an overview of each method’s performance as predictors of free or occupied states in datasets Forest II and III. Results are summarised in Table I, where AUC values closer to one indicate better prediction rates. For increasing levels of perturbation, both approaches experience reduction of AUC, which is anticipated since higher errors on sensor poses negatively affect the quality of occupancy values in both methods. However, in both datasets, the UA-occupancy method displayed higher AUC than the standard occupancy, showing the benefits of fusing pose uncertainty in the occupancy calculation.
@@
TABLE I: Comparison of the AUC values ($\pm$ std. deviation) between the standard occupancy (standard) and UA-occupancy (ours) approaches for increasing perturbation levels in sensor poses, where $\sigma=(0.015\text{m},0.015\text{m},0.015\text{m},0.11^{\circ},0.11^{\circ},0.11^{\circ})$. Higher AUC values indicate better map predictions.
Perturbation Forest II - Trajectory B Forest III - Trajectory C
level standard ours standard ours
[…]
@@ VI-B Full system evaluation
To assess whether our navigation approach (mapping and planning), produces more efficient paths in comparison with a naïve baseline planner without prior maps, we first created a simulated forest with ground obstacles. Then, we simulated two types of aerial sensor trajectory-estimation with perturbed poses, a GPS-only trajectory similar to Sec. VI-A using $\sigma_{\textit{GPS}}=(0.05\text{m},0.05\text{m},0.05\text{m},0.1^{\circ},0.1^{\circ},0.1^{\circ})$, and a SLAM-like trajectory where perturbations (added to relative pose transformations) are accumulated along the trajectory to simulate odometry drift. Absolute corrections (e.g., perturbed GPS observations) are also added at each 120 sec to limit drift. The Gaussian noise added to relative transformations is defined by $\sigma_{\textit{drift}}=(0.009\text{m},0.009\text{m},0.002\text{m},0.0057^{\circ},0.0057^{\circ},0.0057^{\circ})$ and the absolute observation noise defined by $\sigma_{\textit{abs}}=(0.05\text{m},0.05\text{m},0.05\text{m},0.1^{\circ},0.1^{\circ},0.1^{\circ})$. We then processed the lidar scans using our mapping method (Sec. IV) to create the obstruction map used by the D* Lite planner. For the MC integration (Sec. IV-B), we sample poses independently from the GPS-like trajectory, whereas pose-sampling from the SLAM-like trajectory uses ancestral sampling (odometry observations) and independent sampling (absolute observations).
Figure 6: (a) and (b) show the planning results for the scenario with the aerial map created using a GPS-like aerial trajectory. (c) and (d) show the same comparisons for the scenario with map obtained from a simulated aerial SLAM trajectory. In all plots, path ratios below 1 (horizontal line) indicate better results. Suffixes UA and standard in the legends refer to prior maps, UA-occupancy and standard occupancy respectively, used by the navigation method. EXP. COST (5) and EXP. COST (20) refer to expected cost with parameters $C_{\text{obst}}=5$ and 20 respectively. Red markers show mean values including outliers.
@@
Figure 7: The satellite image (a) and the aggregated aerial lidar scans (b) of the forest area used in our experiments. The resulting obstruction map generated using our proposed approach is shown in (c) along with a magnified section for better visualisation in (d).
@@
We defined 40 random start and end goal pairs in the synthetic forest to be executed by our navigation method and the naïve planner. The naïve planner does not use prior maps, and is equivalent to assuming a direct path from the start to the goal as the global prior path. All 40 runs were repeated with three configurations of cost function: log-reachability (LOG-REACH), and two versions of expected cost (EXP. COST) (4) with values $C_{\text{obst}}=5$ and $20$, respectively. During online execution of the planners, we defined a 5 meter radius around the robot pose to update the maps with local information from the ground-truth environment, hence map cells in this region are changed to near 1 or 0, in case obstacles or free areas are detected, respectively, from the ground-truth map. In both methods we use D* Lite to replan in case obstacles block the robot’s planned path. While our method replans on a combined map (prior map with local updates), the naïve method replans only on the map built online with local information.
The metric used for evaluation is the ratio between the executed path length of our navigation method and the naïve baseline. We also compare the executed path length ratio between our planning using UA-occupancy map against planning using standard occupancy, to assess the benefits of our uncertainty incorporation during navigation. Fig. 6 shows the planning results in the two simulated scenarios with maps generated from GPS and SLAM-like aerial trajectories. From Fig. 6 (a), LOG-REACH cost, using UA and standard, both resulted in a significantly higher number of shorter paths than the naïve approach, demonstrating the advantage of using a prior aerial map for global planning and online execution. The EXP. COST, however, showed lower performance with path costs closer to the naïve baseline. This performance reduction is due to some paths planned with expected cost crossing a small number of cells with high obstruction scores instead of planning longer detours, resulting in expensive replans during online execution if the planned path was blocked. Further, although the naïve approach has no prior, in cases with fewer obstructions the naïve assumption (straight-line) may actually be globally optimal or near-optimal, requiring little or no replanning. Because the log-reachability cost is more risk averse by design, it may prefer to detour even small potentially obstructed areas, yielding better results (see Fig. 6 (a)). Simulating GPS-like noise (Fig. 6 (b)) did not show significant performance differences between planning with an uncertainty-aware map versus standard, confirming the intuitive fact that with an accurate aerial trajectory estimate the difference between planning with UA and standard maps is not significant. Planning behaviour differences become more evident as uncertainty increases, as illustrated by simulations with aerial trajectories with higher noise and drift (SLAM-like). In this new scenario (SLAM-like aerial trajectory) (Fig. 6 (d)), the LOG-REACH using UA maps clearly outperformed the standard counterpart, confirming that in scenarios with higher uncertainty in aerial trajectories, our mapping with pose-uncertainty can improve the planned paths. An additional consequence of the increased noise is that all methods showed reduced performance with respect to naïve (see Fig. 6 (c)). This is expected, because as shown in Table I, higher perturbations in the aerial trajectory result in more degraded maps. However, even in this higher noise scenario, using our uncertainty-aware prior aerial map with LOG-REACH is still a better choice than naïve on average (see mean values shown with red markers in Fig. 6). We expect that in more complex terrains, the ground-view sensor’s limited look-ahead (due to significant occlusions, which our current simulation setup does not capture) degrades the performance of naïve even more, ultimately risking the mission completion.
@@ VII Real-World Experiments
Figure 8: Real world experiments Test 1 (a) and Test 2 (b) showing the executed paths of our approach and the naïve method. The start and goal are represented by a black circle and star, respectively. In both experiments our approach guided the robot to shorter, less cluttered paths, whereas the naïve method resulted in a longer trajectory in (a) due to crossing a more cluttered area, and ultimate failure of the planner in (b) by facing a dead-end of high undergrowth.
@@
Since it is not trivial to define ground-truth for obstructed areas in forest environments given their complexity, e.g., cells that are visually “obstructed” by vegetation are often traversable, we qualitatively analysed our mapping output from real data. As observed in Fig. 7 (a)-(d), even with partial canopy occlusion, our mapping using above-canopy lidar scans is capable of detecting regions at ground level with high obstruction scores (yellow to red) indicating the presence of obstacles, and free regions (green) with higher chances of traversing. Uncertain areas (blue) are also captured, providing meaningful information for planning and decision-making.
@@
To test our full system, we defined two start and end goal pairs in the forest named Test 1 and 2. For each test location we performed two runs, one with the naïve approach where the robot used only the local perception and its embedded navigation system to reach the target without any prior information, and another run using our proposed navigation method (prior aerial mapping and global planning) with the log-reachability cost as it demonstrated superior performance in simulation. Unlike the navigation pipeline used in simulations, here our method was used only offline to create the obstruction map and plan a global trajectory from start to goal using D* Lite. For online operation, the planned trajectory was provided to the robot as a set of waypoints (transformed from the aerial map’s reference frame to the robot’s local frame). During online execution, the robot uses the local navigation system [21] with the Hybrid A* planner [28], and onboard perception to track the waypoints while also navigating around locally-detected obstacles. For fair comparison, the same local navigation system was used for both runs, naïve and our method; in this manner, we can assess the benefits of providing information prior to operation through a global trajectory.
@@
During Test 1 (Fig. 8 (a)), while our approach preferred a less cluttered path on the left side of the large obstacle, the naïve approach took the right side, following a much denser path that also crossed a narrow passage among high bushes. Our method travelled a path length of $69.6$ m in comparison to $79.9$ m of the naïve trajectory, which demonstrates that our proposed navigation not only increases path efficiency, but it is also more risk averse preferring safer areas with less likelihood of having obstacles. In Test 2 (Fig. 8 (b)) the limited look-ahead of the naïve approach directed the robot to a dead-end composed of cluttered high vegetation. The local navigation was unable to detour the blocked area, and failed to reach the goal. Unlike the naïve counterpart, our method leveraged the prior information, avoiding the dead-end and taking a clear path on the right of the vegetated area to reach the goal.
@@
Although we acknowledge that the practical experiments are not statistically representative results, the outcomes validate the beneficial behaviours of our method observed in simulation. The field trials showed that planning using our obstruction maps in tandem with the proposed log-reachability cost leads to shorter and safer paths by avoiding cluttered areas, which ultimately increases the mission success rate.
Our mapping module runtime on an Intel Core i7 with 62.4GB RAM (4 CPUs) is approximately $37$ min for 28 million lidar points, sufficient for offline mapping prior to ground robot deployment. For realtime applications, GPU processing could be used to reduce the mapping runtime.